Find the domain of the following function. State your answer in interval notation. Please show all of your work.
f(x)=8x+15/ √22x-121-x

Answer by lwsshak3(11628)

Find the domain of the following function. State your answer in interval notation. Please show all of your work.
f(x)=8x+15/ √22x-121-x
...
Set denominator=0 to find which x-values would make f(x) undefined
√22x-121-x=0
√22x=121+x
square both sides
22x=121^2+2*121x+x^2
22x=14641+242x+x^2
x^2+220x+14641=0
This quadratic equation has a pair of imaginary roots; for this reason, the denominator cannot be equal to zero to make f(x) undefined. However, there is a restriction in the denominator.
√22x
22x≥0
or x≥0
Domain of f(x): [0,∞)